# UNK-VQA: A Dataset and A Probe into Multi-modal Large Models’ Abstention Ability

### v3

@@ VI-C Vision-Language Transformers
Building upon the achievements of large language models (LLMs), researchers have explored the potential of vision-enhanced LLM foundation models. They have extensively investigated the fundamental capabilities of LLMs, such as instruction tuning, in-context learning, and chain-of-thought, for multi-modal large models. Notably, several distinguished models including Flamingo [9], InstructBLIP [15], and Otter [14], have demonstrated the ability to rapidly adapt to new tasks with a few examples.
@@ VII Conclusion and Future Work
This paper proposes a novel dataset to enable VQA models with the ability of abstention. We realize that refraining from answering questions that one does not know is a bedrock of intelligence. Therefore, testing models for this ability is important. By leveraging this dataset, we probe the robustness of multi-modal large models and identify their relatively less satisfactory performance. We then propose a straightforward approach to help train VQA models with the inclusion of unanswerable questions. We evaluate the effectiveness of this approach by integrating it into several baseline methods. The findings reported in this paper highlight two promising future research directions:
1) Building more challenging benchmarks as the VQA problem is yet far from being solved. While large models have shown impressive results on conventional datasets, they often lack robustness when faced with small perturbations, as demonstrated in this paper. To enhance the trustworthiness and generalizability of large models, it is imperative to collect more diverse data and construct more comprehensive benchmarks.
2) Exploring additional possibilities for better multi-modal large model training. The existing VL large model is still in its infancy, with most efforts focused on aligning them with LLMs. However, we have discovered that these large models exhibit significant limitations when it comes to tasks such as instruction following and visual understanding. As such, a more promising direction could involve building multi-modal large models that leverage the intrinsic attributes of multi-modality, rather than relying solely on mapping it back to language.

### v4

@@ VI-C Vision-Language Transformers
Building upon the achievements of large language models (LLMs), researchers have explored the potential of vision-enhanced LLM foundation models. They have extensively investigated the fundamental capabilities of LLMs, such as instruction tuning, in-context learning, and chain-of-thought, for multi-modal large models. Notably, several distinguished models including InstructBLIP [15], MiniGPT-4 [78], and LLaVA [16], have demonstrated the ability to rapidly adapt to new tasks with a few examples.
@@ VII Conclusion and Future Work
This paper proposes a novel UNK-VQA dataset to enable VQA models with the ability of abstention. We realize that refraining from answering questions that one does not know is a bedrock of intelligence. Therefore, testing models for this ability is important. By leveraging this dataset, we probe the robustness of multi-modal large models and identify their relatively less satisfactory performance. We then propose a straightforward approach to help train VQA models with the inclusion of unanswerable questions. We evaluate the effectiveness of this approach by integrating it into several baseline methods. The findings reported in this paper highlight two promising future research directions:
I) Building more challenging benchmarks as the VQA problem is yet far from being solved. While large models have shown impressive results on conventional datasets, they often lack robustness when faced with small perturbations, as shown in this paper. To enhance the trustworthiness and generalizability of large models, it is imperative to collect more diverse data and construct more comprehensive benchmarks.
II) Exploring additional possibilities for better multi-modal large model training. The existing VL large model is still in its infancy, with most efforts focused on aligning them with LLMs. However, we have discovered that these large models exhibit significant limitations on tasks such as instruction following and visual understanding. As such, a more promising direction could involve building multi-modal large models that leverage the intrinsic attributes of multi-modality, rather than relying solely on mapping it back to language.